OpenStudy (anonymous):

A curve y = f(x) has the property that the slope of tangent to the curve at every point is reciprocal to the y value of the curve. Write down a differential equation whose solution is y = f(x). No need to solve this equation.

zepdrix (zepdrix):

If our function is y... `slope of tangent to the curve at every point` that would be this,\[\Large\bf\sf \frac{dy}{dx}\] `is`\[\Large\bf\sf =\] `the reciprocal to the y value of the curve`\[\Large\bf\sf \frac{1}{y}\]

zepdrix (zepdrix):

\[\Large\bf\sf \frac{dy}{dx}\quad=\quad \frac{1}{y}\]Mmm I think that's all they were looking for, yes? :o

OpenStudy (anonymous):

yea thats it thanks